OpenStudy (anonymous):

If AC = 36 and BC = 18, find AD. Round to the nearest tenth, if necessary.

jimthompson5910 (jim_thompson5910):

using AC and BC, what is AB?

OpenStudy (anonymous):

i dont know, 18? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you familiar with the pythagorean theorem?

OpenStudy (anonymous):

sort of

jimthompson5910 (jim_thompson5910):

it's the formula a^2 + b^2 = c^2 where 'a' and 'b' are the legs of a right triangle, c is the hypotenuse 

jimthompson5910 (jim_thompson5910):

compare the triangle I just drew to the triangle you have for your problem

jimthompson5910 (jim_thompson5910):

you'll see that a = unknown b = 18 c = 36

OpenStudy (anonymous):

18 square root 3?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

now it turns out that there are 3 triangles in your picture they are: triangle ABC, which is the biggest triangle (B is the right angle) triangle ADB, where D is the right angle triangle BDC, where D is the right angle we'll focus on triangles ABC and ADB why? Because we know AB, AC and BC, it makes sense to use ABC. We want to find AD, so that's why we're using triangle ADB

jimthompson5910 (jim_thompson5910):

If we focus on triangle ABC, we see that AC is the hypotenuse If we turn our attention to triangle ADB, we'll find that AB is the hypotenuse So they are corresponding parts

jimthompson5910 (jim_thompson5910):

AD is the longer leg of triangle ADB AB is the longer leg of triangle ABC so this is another pair of corresponding parts

jimthompson5910 (jim_thompson5910):

we then divide the corresponding parts of ABC over ADB (in the same order) like this AC/AB = AB/AD then plug in the info we know 36/(18*sqrt(3)) = (18*sqrt(3))/AD and then solve for AD

OpenStudy (anonymous):

and what would be the answer?

jimthompson5910 (jim_thompson5910):

were you able to solve for AD?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

cross multiply to get 36*AD = (18*sqrt(3))(18*sqrt(3)) 36*AD = 18*18*sqrt(3)*sqrt(3) 36*AD = 18^2*(sqrt(3))^2 36*AD = 324*3 36*AD = 972 I'll let you finish

OpenStudy (anonymous):

i dont understand what you are doing and what sqrt is

jimthompson5910 (jim_thompson5910):

sqrt stands for "square root"

OpenStudy (anonymous):

okay and what would the answer be?

jimthompson5910 (jim_thompson5910):

look back at my steps and finish up solving for AD

OpenStudy (anonymous):

@jim_thompson5910 im on this question right now can you help me? pls